# The role of charm and unflavored mesons in prompt atmospheric lepton fluxes

### v1

@@ 5 Discussion and Conclusions
Figure 16: Left: Angle-averaged (up-to $\theta_{\rm zen}\leq 60^{\circ}$) electron neutrino flux ($\nu_{e}+\bar{\nu}_{e}$) scaled by $E_{\nu_{e}}^{3}$ with the inclusion of intrinsic charm with $w_{c}^{\rm intr}$ values of $3.87\times 10^{-3}$ (dashed lines) and $4.65\times 10^{-4}$ (solid lines). Right: Angle-averaged for full zenith angle range, the $\nu_{e}+\bar{\nu}_{e}$ flux contributions scaled by $E_{\nu_{e}}^{2}$, along with a data point from IceCube IceCube:2015mgt .
We have used IceCube’s upper bound on the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ flux from measurements of through-going muons IceCube:2016umi to constrain the intrinsic charm contribution. IceCube’s upper bound relies on the Enberg et al. (ERS) Enberg:2008te prompt neutrino flux, modified by an input H3p cosmic ray flux and a single power law model of the diffuse astrophysical neutrino flux. A recent re-evaluation of the IceCube sensitivity from muon tracks, using H4a and GST all-nucleon cosmic ray fluxes and a single power law for the astrophysical flux, yields an 90% CL upper bound on the prompt flux normalization of 2.9 times the Sibyll-2.3c prompt neutrino flux IceCube:2023mrq . With either cascades or northern tracks and a single power law astrophysical flux, the 90% CL upper bound on the prompt flux normalization is 1.34 and 3.02 times the Sibyll-2.3c prompt neutrino flux, respectively Bottcher:2025hcz . Upper limits vary with the assumed astrophysical neutrino flux, and whether or not combined fits are made. Despite the model dependence of the upper bound on the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ flux, we conclude that there is room for only a small intrinsic charm contribution to the prompt neutrino flux.
@@
Measurements of the high-energy atmospheric $\nu_{e}+\bar{\nu}_{e}$ flux may shed light on the prompt flux contribution. While the prompt atmospheric flux of $\nu_{e}+\bar{\nu}_{e}$ is very nearly equal to the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ flux, the conventional fluxes have larger differences. The conventional atmospheric $\nu_{e}+\bar{\nu}_{e}$ flux is about a factor of 20 lower than the conventional $\nu_{\mu}+\bar{\nu}_{\mu}$ flux because nearly 100% decays of $\pi^{+}$ lead to $\nu_{\mu}+\mu^{+}$, and at the energies under consideration here, the muons do not decay. Kaon production and decay dominate as the source of conventional atmospheric $\nu_{e}+\bar{\nu}_{e}$ flux. Consequently, the prompt $\nu_{e}+\bar{\nu}_{e}$ flux becomes larger than the corresponding conventional flux at energies lower than in the case of $\nu_{\mu}+\bar{\nu}_{\mu}$. This is illustrated in fig. 16, where we show both pQCD and two versions of the prompt flux, both with and without including intrinsic charm. In the left panel of fig. 16, the $\nu_{e}+\bar{\nu}_{e}$ flux is scaled by $E_{\nu_{e}}^{3}$ (as with the prior flux figures). As already noted, the dashed lines that include $w_{c}^{\rm intr}=3.87\times 10^{-3}$ violate the upper limit for the prompt $\nu_{\mu}+\bar{\nu}_{\mu}$ flux.
@@
The angle-averaged atmospheric electron neutrino flux was measured by DeepCore/IceCube for $E_{\nu}=80$ GeV $\penalty 10000\ -\penalty 10000\ 6$ TeV IceCube:2012jwm and up to a highest energy bin at $\log_{10}(E_{\nu}/{\rm GeV})=4.0-5.0$ by IceCube IceCube:2015mgt . The right panel of fig. 16 shows the fluxes in the left panel, now scaled by $E_{\nu_{e}}^{2}$, and the data point of the highest energy bin of the IceCube atmospheric electron neutrino flux measurement IceCube:2015mgt . The measurements of the atmospheric $\nu_{e}+\bar{\nu}_{e}$ flux is difficult at high energy, in part because the diffuse astrophysical neutrino flux is large. For reference, the IceCube’s high-energy starting event measurements of the per-flavor astrophysical flux in the energy region of $E_{\nu}=100$ TeV, assuming a single power law, is $E_{\nu}^{2}\phi_{\nu+\bar{\nu}}^{\rm per\ flavor}\sim 2\times 10^{-8}(E_{\nu}/100\ {\rm TeV})^{-2.5}$ IceCube:2020acn ; IceCube:2024fxo ; IceCube:2025dlr . To the extent that the diffuse astrophysical neutrino flux is inferred from $\nu_{\mu}+\bar{\nu}_{\mu}$ measurements, it may be possible to understand the relative contributions of the atmospheric and prompt $\nu_{e}+\bar{\nu}_{e}$ fluxes in the measurements.
@@
Figure 17: Angle-averaged ($\theta_{\rm zen}\leq 60^{\circ}$) atmospheric $\phi_{\mu^{+}+\mu^{-}}$ scaled by $E_{\mu}^{3.7}$ with $w_{c}^{\rm intr}=0$ and $\alpha=1,\ 3.81,\ 4.80,\ 5.78$ and $8.60$. The IceCube data from ref. IceCube:2015wro are also shown.
The IceCube limits on the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ flux put significant constraints on the possibility that intrinsic charm can account for IceCube’s high energy $\mu^{+}+\mu^{-}$ measurements, albeit with large errorbars. The measurements suggest a prompt component that could otherwise come from prompt decays of light unflavored mesons, as we discussed in section 4 with an enhancement factor of $\alpha_{\rm unfl}=4.80$ larger than the MCEq default contribution. The IceCube Collaboration reports that the combined heavy-flavor prompt plus unflavored prompt contributions to the atmospheric muon flux correspond to 4.75 times the ERS flux with a 90% CL range of 2.33–9.34 times the ERS flux, assuming the H3a all-nucleon cosmic ray energy spectrum IceCube:2015wro . Their conclusion is consistent with their angular analysis as well. IceCube’s results are consistent with the enhancements of the prompt contributions to the muon flux discussed here.
@@
We again show the large experimental errors in the measured $\mu^{+}+\mu^{-}$ flux scaled by $E_{\mu}^{3.7}$ in fig. 17, where the experimental data are compared to theory predictions for different values of $\alpha_{\rm unfl}$ and $w^{c}_{\rm intr}=0$. In particular, the figure shows curves for $\alpha_{\rm unfl}=1$, and $\alpha_{\rm unfl}=4.8^{+0.98}_{-0.99}$, corresponding to the $\chi^{2}$ best-fit value with $1\sigma$ errors. We also show a curve for $\alpha=8.6$, to show a more extreme scaling. It is still consistent within $2\sigma$ uncertainties. All of the scaling factors multiply the unflavored meson contributions to the prompt flux according to interactions modeled by Sibyll-2.3c. As the lower panel of fig. 13 shows, DPMJETIII306 would lead to a larger high-energy contribution. However, its energy dependence cannot account for the apparent excess of atmospheric muons in the hundreds of TeV energy range.
@@
The extraction of the prompt atmospheric muon and neutrino fluxes from neutrino telescope data remains a topic of tremendous interest. In combination, SHiP Alekhin:2015byh ; Bai:2018xum and the forward physics program Anchordoqui:2021ghd ; Feng:2022inv with neutrinos Bai:2022xad ; Bai:2022jcs and studies of unflavored meson production at accelerator experiments, in combination with a better modeling of the unflavored component in Monte Carlo event generators for (astro)-particle physics and more precise tunings of the relevant parameters of the latter Albrecht:2025kbb , will be important to better interpret existing and future data, potentially resolving the tension between current theoretical predictions and measurements of the high-energy atmospheric $\mu^{+}+\mu^{-}$ flux.
@@ Acknowledgements.
We are grateful to Anatoli Fedynitch for help and tips on the use of the MCEq framework, to Felix Riehn for discussions on the physics implemented in the Sybill event generator and comparisons with our QCD predictions and to Sergey Ostapchenko and Johannes Blümlein for sharing with us their considerations on intrinsic charm. The model options for intrinsic charm used in this work were suggested by Sergey Ostapchenko. This work was supported in part by the US Department of Energy under grant DE-SC-0010113, by the Deutsche Forschung Gemeinschaft within the FOR2926 Resarch Unit (Project ID 40824754) and under Germany’s Excellence Strategy – EXC 2121 “Quantum Universe” – 390833306. For facilitating portions of this research, DG wishes to acknowledge the Center for Theoretical Underground Physics and Related Areas (CETUP*), The Institute for Underground Science at Sanford Underground Research Facility (SURF), and the South Dakota Science and Technology Authority for hospitality and financial support, as well as for providing a stimulating environment.
@@
Traditionally, the $\mu^{+}$ energy distribution from $\Lambda_{c}$ decays is evaluated using three decay form factors Buras:1976dg ; Pietschmann:1984en . (See also the appendix of the preprint version of ref. Bugaev:1998bi .) We evaluated the energy distributions for $\mu^{+}$ and $\nu_{\mu}$ in an effective three-body semi-leptonic decay including the muon mass and an effective final-state strange hadron mass $m_{s}^{\rm eff}=1.27$ GeV Bugaev:1998bi . The $\mu^{+}$ and $\nu_{\mu}$ $x$-distribution in the rest frame and $y$-distribution in the boosted frame are shown in the left and right panel of fig. 19.
It was noted already in Ref. Illana:2010gh ; Volkova:2011zza that, due to the VAV-A couplings of leptons in the weak charged current, neutrinos carry about 15% more of the parent charm hadrons’ momentum than the muon. We confirm this with our distributions in the $Z$-moment approximation. In this approximation, the lepton flux is proportional to $Z_{h\ell}$, for hadron $h$ and lepton $\ell=\mu,\nu$, where
@@
in terms of the lepton energy $E$, hadron energy $E_{h}$, the flux of hadrons $\phi_{h}$ that depends on energy and the distance traveled in a mean $h$ decay time $d_{h}(E)\propto E$. For an incident cosmic ray flux that has power law scaling $\phi_{\rm CR}\propto E^{-\gamma}$, charm hadrons in the prompt regime scale as $\phi_{h}\propto E^{-\gamma+1}$ Lipari:1993hd ; Gondolo:1995fq . The decay $Z$-moment is energy independent in this approximation and can be written as

### v2

@@ 5 Discussion
Figure 17: Left: Angle-averaged (up-to $\theta_{\rm zen}\leq 60^{\circ}$) electron neutrino flux ($\nu_{e}+\bar{\nu}_{e}$) scaled by $E_{\nu_{e}}^{3}$ with the inclusion of intrinsic charm with $w_{c}^{\rm intr}$ values of $3.93\times 10^{-3}$ (dashed lines) and $1.01\times 10^{-3}$ (solid lines). Right: Angle-averaged for full zenith angle range, the $\nu_{e}+\bar{\nu}_{e}$ flux contributions scaled by $E_{\nu_{e}}^{2}$, along with a data point from IceCube Aartsen and others (2015).
@@
Measurements of the high-energy atmospheric $\nu_{e}+\bar{\nu}_{e}$ flux may shed light on the prompt flux contribution. While the prompt atmospheric flux of $\nu_{e}+\bar{\nu}_{e}$ is very nearly equal to the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ flux, the conventional fluxes have larger differences. The conventional atmospheric $\nu_{e}+\bar{\nu}_{e}$ flux is about a factor of 20 lower than the conventional $\nu_{\mu}+\bar{\nu}_{\mu}$ flux because nearly 100% decays of $\pi^{+}$ lead to $\nu_{\mu}+\mu^{+}$, and at the energies under consideration here, the muons do not decay. Kaon production and decay dominate as the source of conventional atmospheric $\nu_{e}+\bar{\nu}_{e}$ flux. Consequently, the prompt $\nu_{e}+\bar{\nu}_{e}$ flux becomes larger than the corresponding conventional flux at energies lower than in the case of $\nu_{\mu}+\bar{\nu}_{\mu}$. This is illustrated in fig. 17, where we show both pQCD and two versions of the prompt flux, both with and without including intrinsic charm. In the left panel of fig. 17, the $\nu_{e}+\bar{\nu}_{e}$ flux is scaled by $E_{\nu_{e}}^{3}$ (as with the prior flux figures). As already noted, the dashed lines that include $w_{c}^{\rm intr}=3.93\times 10^{-3}$ violate the upper limit for the prompt $\nu_{\mu}+\bar{\nu}_{\mu}$ flux.
@@
The angle-averaged atmospheric electron neutrino flux was measured by DeepCore/IceCube for $E_{\nu}=80$ GeV $\penalty 10000\ -\penalty 10000\ 6$ TeV Aartsen and others (2013b) and up to a highest energy bin at $\log_{10}(E_{\nu}/{\rm GeV})=4.0-5.0$ by IceCube Aartsen and others (2015). The right panel of fig. 17 shows the fluxes in the left panel, now scaled by $E_{\nu_{e}}^{2}$, and the data point of the highest energy bin of the IceCube atmospheric electron neutrino flux measurement Aartsen and others (2015). The measurements of the atmospheric $\nu_{e}+\bar{\nu}_{e}$ flux is difficult at high energy, in part because the diffuse astrophysical neutrino flux is large. For reference, the IceCube’s high-energy starting event measurements of the per-flavor astrophysical flux in the energy region of $E_{\nu}=100$ TeV, assuming a single power law, is $E_{\nu}^{2}\phi_{\nu+\bar{\nu}}^{\rm per\ flavor}\sim 2\times 10^{-8}(E_{\nu}/100\ {\rm TeV})^{-2.5}$ Aartsen and others (2020); Abbasi and others (2024); Balagopal V. and others (2025). To the extent that the diffuse astrophysical neutrino flux is inferred from $\nu_{\mu}+\bar{\nu}_{\mu}$ measurements, it may be possible to understand the relative contributions of the atmospheric and prompt $\nu_{e}+\bar{\nu}_{e}$ fluxes in the measurements.
@@
Figure 18: Angle-averaged ($\theta_{\rm zen}\leq 60^{\circ}$) atmospheric $\phi_{\mu^{+}+\mu^{-}}$ scaled by $E_{\mu}^{3.7}$ with $w_{c}^{\rm intr}=0$ and $\alpha=1,\ 3.87,\ 4.86,\ 5.85$ and $8.72$. The IceCube data from ref. Aartsen and others (2016a) are also shown.
@@
We again show the large experimental errors in the measured $\mu^{+}+\mu^{-}$ flux scaled by $E_{\mu}^{3.7}$ in fig. 18, where the experimental data are compared to theory predictions for different values of $\alpha_{\rm unfl}$ and $w^{c}_{\rm intr}=0$. In particular, the figure shows curves for $\alpha_{\rm unfl}=1$, and $\alpha_{\rm unfl}=4.86^{+0.99}_{-0.99}$, corresponding to the $\chi^{2}$ best-fit value with $1\sigma$ errors. We also show a curve for $\alpha=8.72$, to show a more extreme scaling. It is still consistent within $2\sigma$ uncertainties. All of the scaling factors multiply the unflavored meson contributions to the prompt flux according to interactions modeled by Sibyll-2.3c. As the lower panel of fig. 14 shows, DPMJETIII306 would lead to a larger high-energy contribution. However, its energy dependence cannot account for the apparent excess of atmospheric muons in the hundreds of TeV energy range.
6 Conclusions
IceCube’s measurements of the $\mu^{+}+\mu^{-}$ atmospheric flux at high energies suggest a missing prompt component. We have focused on fluxes at zenith angles $\theta\leq 60^{\circ}$ to avoid issues with potential mis-reconstruction of muon events in the experimental data Aartsen and others (2016a). Since the prompt $\nu_{\mu}+\bar{\nu}_{\mu}$ and $\mu^{+}+\mu^{-}$ fluxes from charm are nearly equal, an increase of the prompt $\mu^{+}+\mu^{-}$ flux from charm to improve agreement with the high-energy muon data (which would correspond to taking $w_{c}^{\rm intr}=3.93\times 10^{-3}$) can violate the upper bounds on the prompt $\nu_{\mu}+\bar{\nu}_{\mu}$ flux. Indeed, in addition, such a high value of $w_{c}^{\rm intr}$ would lead to an excess of very forward production of $\Lambda_{c}^{+}$ in low-energy collisions since we have assumed that $w_{c}^{\rm intr}$ is energy independent. The value of $w_{c}^{\rm intr}=1.01\times 10^{-3}$ ($w_{c}^{\rm intr}=5.34\times 10^{-4}$) determined from the 2025 (2016) IceCube prompt neutrino limit is sufficiently small that the intrinsic charm cross section is less than $\sim 7\%$ (4%) of the perturbative cross section for fixed-target $pp$ production of $c\bar{c}$ at $E_{p}=10^{5}$ GeV and less than $\sim 3\%$ (2%) for $E_{p}\gtrsim 10^{6}$ GeV, well within the next-to-leading order perturbative Nason et al. (1988); Bhattacharya et al. (2015, 2016); Garzelli et al. (2015) and measurement Adare and others (2006); Abt and others (2007); Abelev and others (2012); Adamczyk and others (2012); Aad and others (2016); Aaij and others (2013) uncertainties for open charm production. However, this smaller value of $w_{c}^{\rm intr}$ still underpredicts the atmospheric $\mu^{+}+\mu^{-}$ flux at high energy, as shown in fig. 7.
We have used IceCube’s upper bound on the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ flux from measurements of through-going muons Abbasi and others (2025a) to constrain the intrinsic charm contribution. This IceCube’s upper bound relies on a Sibyll-2.3c-based MCEq evaluation of charm production and an input H3a cosmic ray flux. With either cascades or northern tracks and a single power law astrophysical flux, the 90% CL upper bound on the prompt neutrino flux normalization is 1.34 and 3.02 times the Sibyll-2.3c prompt neutrino flux, respectively Böttcher (2025). Upper limits vary with the assumed astrophysical neutrino flux, with normalization factors ranging between 1.92-2.99 Abbasi and others (2025a), and whether or not combined fits are made. Despite the model dependence of the upper bound on the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ flux, we conclude that there is room for only a small intrinsic charm contribution to the prompt neutrino flux.
The IceCube limits on the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ flux put significant constraints on the possibility that intrinsic charm can account for IceCube’s high energy $\mu^{+}+\mu^{-}$ measurements, albeit with large errorbars. The measurements suggest a prompt component that could otherwise come from prompt decays of light unflavored mesons, as we discussed in section 4 with an enhancement factor of $\alpha_{\rm unfl}=4.86$ larger than the MCEq default contribution. The IceCube Collaboration reports that the combined heavy-flavor prompt plus unflavored prompt contributions to the atmospheric muon flux correspond to 4.75 times the ERS flux with a 90% CL range of 2.33–9.34 times the ERS flux, assuming the H3a all-nucleon cosmic ray energy spectrum Aartsen and others (2016a). Their conclusion is consistent with their angular analysis as well. IceCube’s results are consistent with the enhancements of the prompt contributions to the muon flux discussed here.
@@ 6 Conclusions
The extraction of the prompt atmospheric muon and neutrino fluxes from neutrino telescope data remains a topic of tremendous interest. In combination, SHiP Alekhin and others (2016); Bai and Reno (2019) and the forward physics program Anchordoqui and others (2022); Feng and others (2023) with neutrinos Bai et al. (2023, 2022) and studies of unflavored meson production at accelerator experiments, together with a better modeling of the unflavored component in Monte Carlo event generators for (astro)-particle physics and more precise tunings of the relevant parameters of the latter Albrecht and others (2025), will be important to better interpret existing and future data, potentially resolving the tension between current theoretical predictions and measurements of the high-energy atmospheric $\mu^{+}+\mu^{-}$ flux. Furthermore, refined results on the cosmic ray spectrum and composition, e.g., as recently reported by the LHAASO collaboration Lv et al. (2025), may have an impact on predictions for prompt atmospheric leptons.
Note Added
This version of the manuscript differs from the version first appearing in arXiv in December 2025 because it includes the new IceCube limit on prompt neutrino fluxes Abbasi and others (2025a) that appeared in December 2025, almost simultaneously to our work. Although the new experimental results have some impact on part of the numerical values extracted and discussed in this work, the conclusions and main message remain essentially unchanged.
@@ Acknowledgements.
We are grateful to Anatoli Fedynitch for help and tips on the use of the MCEq framework, to Felix Riehn for discussions on the physics implemented in the Sibyll event generator and comparisons with our QCD predictions and to Sergey Ostapchenko and Johannes Blümlein for sharing with us their considerations on intrinsic charm. The model options for intrinsic charm used in this work were suggested by Sergey Ostapchenko. This work was supported in part by the US Department of Energy under grant DE-SC-0010113, by the Deutsche Forschung Gemeinschaft within the FOR2926 Resarch Unit (Project ID 40824754) and under Germany’s Excellence Strategy – EXC 2121 “Quantum Universe” – 390833306. For facilitating portions of this research, DG wishes to acknowledge the Center for Theoretical Underground Physics and Related Areas (CETUP*), The Institute for Underground Science at Sanford Underground Research Facility (SURF), and the South Dakota Science and Technology Authority for hospitality and financial support, as well as for providing a stimulating environment.
@@
Traditionally, the $\mu^{+}$ energy distribution from $\Lambda_{c}$ decays is evaluated using three decay form factors Buras (1976); Pietschmann (1984). (See also the appendix of the preprint version of ref. Bugaev et al. (1998).) We evaluated the energy distributions for $\mu^{+}$ and $\nu_{\mu}$ in an effective three-body semi-leptonic decay including the muon mass and an effective final-state strange hadron mass $m_{s}^{\rm eff}=1.27$ GeV Bugaev et al. (1998). The $\mu^{+}$ and $\nu_{\mu}$ $x$-distribution in the rest frame and $y$-distribution in the boosted frame are shown in the left and right panel of fig. 20.
It was noted already in Ref. Illana et al. (2011); Volkova (2011) that, due to the VAV-A couplings of leptons in the weak charged current, neutrinos carry about 15% more of the parent charm hadrons’ momentum than the muon. We confirm this with our distributions in the $Z$-moment approximation. In this approximation, the lepton flux is proportional to $Z_{h\ell}$, for hadron $h$ and lepton $\ell=\mu,\nu$, where
@@
in terms of the lepton energy $E$, hadron energy $E_{h}$, the flux of hadrons $\phi_{h}$ that depends on energy and the distance traveled in a mean $h$ decay time $d_{h}(E)\propto E$. For an incident cosmic ray flux that has power law scaling $\phi_{\rm CR}\propto E^{-\gamma}$, charm hadrons in the prompt regime scale as $\phi_{h}\propto E^{-\gamma+1}$ Lipari (1993); Gondolo et al. (1996). The decay $Z$-moment is energy independent in this approximation and can be written as